OpenStudy (anonymous):

Preston began solving the following equation. His work so far is shown. What should Preston do next? k divided by four plus one-eighth equals seven-eighths. Subtract one-eighth from both sides: k divided by four plus one-eighth minus one-eighth equals seven-eighths minus one-eighth. So, k divided by four equals six-eighths.

OpenStudy (adamk):

k/4 + 1/8 = 7/8 -- subtract 1/8 k/4 = 7/8 - 1/8 -- combine like terms k/4 = 6/8 -- MULTIPLY BOTH SIDES BY 4 k = 6/8(4) k = 24/8 k = 3